Fermi National Laboratory
 

Operation Summary for January 22, 2010 - January 25, 2010

 

The Integrated Luminosity for the period from 1/18/10 to 1/25/10 was 47.72 pb-1.

NuMI reported receiving 6.96E18 protons on target during this same period.

Friday Jan. 22

The day shift began with Operations monitoring the stash (329.4E10), with the Antiproton Source stacking (22.3E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest experiments taking beam, with the Collider Detector at Fermilab (CDF) in access, and with the D-Zero (D0) experiment waiting for the next store.

During the day shift, CDF came out of access at 10 AM. TeV personnel conducted orbit studies until 3 PM, which ended with a TeV quench at sector F4 due to abort kicker problems. Quench recovery lasted through the rest of the shift.

During the evening shift, the TeV sector F4 cryo system recovered at 7:30 PM, but problems with a B0 power supply readback held off HEP shot setup until 11 PM. Operations disabled the MTest beam at 6 PM.

Saturday Jan. 23

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (118.6E10), with Pbar stacking (100.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 7530 at 12:45 AM with an average initial luminosity (AIL) of 290.85E30. Operations established beam to MTest at 4:30 AM.

During the day shift, A Linac Klystron RF station high voltage cable faulted early in the morning and held off beam to all users until 12:45 PM. Operations terminated store 7530 at 3:15 PM.

During the evening shift, Operations prepared D0 for a collision hall access at 4:45 PM. After the access, Operations conducted TeV orbit checks. Operations put the TeV into shot setup and established store 7533 at 8 PM with an AIL of 266.9E30. At 11 PM, a Booster capacitor in the tunnel failed and held off beam. Booster personnel accessed to investigate.

Sunday Jan. 24

The midnight shift began with the Booster in access, with Operations monitoring store 7533, stash, and stack, and with NuMI and MiniBooNE off due to the Booster access.

During the midnight shift, Booster personnel had the capacitor bypassed by 12:45 AM. At 4:30 AM, a Booster vacuum gauge failure caused another hour of downtime to users. Operations established beam to MTest at 4 AM. Operations reported that the H- Source output had dropped to 33 mA.

During the day shift, Operations terminated store 7533 at 2:30 PM and then put the TeV into shot setup. Operations reported that the Proton Source output continued to decline through the day, which limited beam to all users.

During the evening shift, Operations established store 7534 at 4:15 PM with an AIL of 236.1E30. Operations disabled beam to MTest at 7 PM.

Monday Jan. 25

The midnight shift began with Operations monitoring store 7534 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4 AM. Beam out of the Proton Source continued to decline through the night, severely limiting beam to users.

The Plans for Monday

The plans for today are to establish a new store as soon as the Proton Source can deliver enough protons and then, during the Tuesday midnight shift, Operations will established another store. The Run Coordinator will place the accelerator complex into a maintenance period for three to four days starting on Tuesday morning while Linac personnel repair or replace the H-Source.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory